1. Bill's age is a year less than twice Jane's age. Five years ago, Bill's age was three times Jane's age. Find Bill's age.

Found 2 solutions by josgarithmetic, ikleyn:
Answer by josgarithmetic(39620)

If necessary, assign variables to the age of each, write the description as equations, and solve the equations. Plenty of these kind of age problems have been solved on the site. This is one of the simpler of them.


Jane, 9
Bill, 17

Answer by ikleyn(52803)

.
B = 2J - 1       (1)    ("Bill's age is a year less than twice Jane's age.")

B-5 = 3*(J-5)    (2)    ("Five years ago, Bill's age was three times Jane's age.")


Substitute expression (1) for B into eq(2), replacing B. You will get


(2J-1) - 5 = 3*(J-5),

2J - 6 = 3J - 15,

15 - 6 = 3J - 2J  ====>  J = 9.


Then  B = 2J-1 = 2*9-1 = 17.


Answer.  Bill is 17 years old.

Solved.

----------------
There is a bunch of lessons on age word problems
    - Age problems and their solutions
    - A fresh formulation of a traditional age problem
    - Really intricate age word problem
    - Selected age word problems from the archive
    - Age problems for mental solution
in this site.

Read them and become an expert in solving age problems.

Also, you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lessons are the part of this online textbook under the topic "Age word problems".


Save the link to this online textbook together with its description

Free of charge online textbook in ALGEBRA-I
https://www.algebra.com/algebra/homework/quadratic/lessons/ALGEBRA-I-YOUR-ONLINE-TEXTBOOK.lesson

to your archive and use it when it is needed.